OpenStudy (anonymous):

Prove the statement using ɛ, ɓ definition of a limit. Lim (1+(1/3)x)=2 x->3

OpenStudy (anonymous):

definition in this case would be given any \(\epsilon>0\) there is a \(\delta\) which you write as a function of \(\epsilon\) that if \(|x-3|<\delta\) then \(|1+\frac{1}{3}x-2|<\epsilon\)

OpenStudy (anonymous):

work backwards \[|\frac{1}{3}x-1|<\epsilon\]\[-\epsilon<\frac{1}{3}x-1<\epsilon\]\[1-\epsilon<\frac{1}{3}x<\epsilon+1\] \[3-3\epsilon<x<3\epsilon+3\]

OpenStudy (anonymous):

the last one is the same as \[|x-3|<3\epsilon\] to you can say "let \(\delta=3\epsilon\)" work the algebra backwards and you have your proof

OpenStudy (anonymous):

not really a coincident that your line with slope \(m\) requires a \(\delta\) of \(\frac{1}{m}\)

OpenStudy (anonymous):

Is there any easy way to do/explain this cause I don't really understand the steps that you are doing.

OpenStudy (zzr0ck3r):

well ok we need to show that if |x-3| < delta then |f(x)-L|<epsilon for any given epsilon so he worked backwards which is always a great way to start these, and started where we need to end up, he showed that if that statement is true, then |x-3|<3epsilon then the proof is done by working backwards we got to |x-3|<3e must hold, and that is nice because that is exactly what we need to assume so assume let epsilon be given if delta = 3epsilopn then\[|x-3|<\delta=3\epsilon\\implies\\|\frac{1}{3}x-1|=\frac{|x-3|}{|3|}\le\epsilon\]

OpenStudy (zzr0ck3r):

i dont think I made anything clearer but I gave it a shot

OpenStudy (anonymous):

Just wondering where he got the \[|x-3| < \delta\]

OpenStudy (zzr0ck3r):

\[|\frac{1}{3}x-1|<\epsilon\\|\frac{x-3}{3}|<\epsilon\\\frac{|x-3|}{|3|}<|x-3|<\epsilon\]

OpenStudy (zzr0ck3r):

\[so\space given\space \epsilon\\let\space\delta=\epsilon\\if\space|x-3|<\delta\\then\\|1+\frac{1}{3}x-2|<\frac{|x-3|}{3}<|x-3| <\epsilon\]